OpenStudy (anonymous):

A regular square pyramid has base edge 5 and height 9. A similar pyramid has height 12. Find the ratio of the lateral areas.

OpenStudy (anonymous):

so.. know the formula--> A = (1/2)bs s2/s1 = 12/9 = 4/3 b2/b1 = 4/3 A2/A1 = 16/9.....hope it helps a bit.

OpenStudy (anonymous):

Just try to remember this regular square pyramid' has an easy formula: LatArea = (1/2) * (Perimeter of Base) * (slant height)